OpenStudy (anonymous):

is secxtanx(1-sin^2x)=sinx a trigonometric identity?

OpenStudy (welshfella):

Hint: 1 - 2sin^2 x = cos^2 x - also convert secx tanx to sin's and cos's and see what you get

OpenStudy (welshfella):

1 - sin^2 x = cos ^2 x sec x = 1/ cos x tan x = sin x / cos x replace these in your identity and see what you can simplify - maybe it will simplify to sin x

OpenStudy (welshfella):

if it does then its a trigonometric identity

OpenStudy (anonymous):

okay i'll try that thank you

OpenStudy (welshfella):

yw

OpenStudy (anonymous):

would the answer be no ?

OpenStudy (welshfella):